OpenStudy (anonymous):

y=3/5(x)-12 find the x and y intercepts

OpenStudy (anonymous):

y intercept: when x = 0 x intercept solve 0 = 3/5x - 12

OpenStudy (anonymous):

huh wait why is y 0

OpenStudy (anonymous):

thus y-intercept = -12

OpenStudy (anonymous):

yes but how do you find x

OpenStudy (anonymous):

x intercepts y=0 0=3/5(x)-12 12*5/3=x 20=x y intercept x=0 y=-12

OpenStudy (anonymous):

0 = 3/5x - 12 3/5 x = 12 x = 12 / 3/5 = 12 * 5/3 = 20 x-intercept = 20

OpenStudy (anonymous):

ohhhh i see how

OpenStudy (anonymous):

medal now